Math

The population of Kenya reached up to 25,000,000 people in 1990. When will it reach 50,000,000 people? Assume an annual rate of increase of 4.1%?
Answer #1

Ao(1 + r) = A(t)

25,000,000(1 + 0.041)t = 50,000,000

1.041t = 2

t = log 2 / log 1.041

t = 17.25...= 17 years 1990 + 17 =

 2007


answered by: Tina Gao